OpenStudy (anonymous):

Evaluate x(-2) for x = -3. (-2) is an exponent! :)

OpenStudy (anonymous):

So x^(-2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/9

OpenStudy (anonymous):

Whenever anything is to the power of a negative sign, that indicates that you must flip the numerator to the denominator.

OpenStudy (anonymous):

So then it would be 1/(x^-2)=1/((-3)^2)=1/9

OpenStudy (anonymous):

*1/(x^2)

OpenStudy (anonymous):

it becomes 1/(x^2) and then you plug in the -3 so its 1/9

OpenStudy (anonymous):

thankkks :)

OpenStudy (anonymous):

Did it make sense? Also it is good to note that sometimes you may have problems that look like this: 1/(x^-2)

OpenStudy (anonymous):

which means you need to switch it from the denominator to the numerator. So it would be x^2. Just a helpful note.